OpenStudy (kpop4life123):

Which of the following is equivalent to (mn)^3? 3mn mn^3 m^3n^3 m^3n

OpenStudy (whpalmer4):

\[(mn)^3 = (mn)*(mn)*(mn) = \]

OpenStudy (kpop4life123):

is it mn^3 ?

jimthompson5910 (jim_thompson5910):

notice how @whpalmer4 wrote out `(mn)*(mn)*(mn)` 3 copies of (mn) multiplied

jimthompson5910 (jim_thompson5910):

gather the m's and multiply them out gather the n's and multiply them out

OpenStudy (kpop4life123):

thats why i chose mn^3

jimthompson5910 (jim_thompson5910):

n*n*n = n^3 is true but you forgot to do the same for the m's

OpenStudy (kpop4life123):

oh then my final answer is m^3n^3

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (kpop4life123):

thanks :)

jimthompson5910 (jim_thompson5910):

np